How i will solve this problem:
Andy's average driving speed for a 4-hours trip was 45 miles per hour. During the first 3 hours he drove 40 miles per hour. What was his average speed for the last hour of his trip?

Answer by mananth(16949)

Average speed = 45 mph
time = 4hours
distance traveled = 4*45 =180 miles.
..
first 3 hours average was 40 mph.
distance traveled = 40*3= 120 miles.
..
Balance 60 miles he drove in 1 hour.
so average speed = 60mph.
..
m.ananth@hotmail.ca